OpenStudy (anonymous):

A strip of wood 78 inches long is to be cut into pieces 3 3/4 inches long. How many pieces can be cut?

OpenStudy (anonymous):

78/(15/4)

OpenStudy (anonymous):

3 3/4=15/4

OpenStudy (anonymous):

or 15/4=3 3/4=3.75

OpenStudy (anonymous):

thus 78/3.75=your answer

OpenStudy (anonymous):

which is.....